OpenStudy (anonymous):

Evaluate the following expression for a = 6, b= 3 and c = 4. a + 2b - c

OpenStudy (anonymous):

20

OpenStudy (anonymous):

lol hii

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

i knew it i was just giving people something to do

OpenStudy (anonymous):

i was wondering why it was easy

OpenStudy (anonymous):

unless i got it wrong

OpenStudy (anonymous):

lol its 8

OpenStudy (anonymous):

woops

OpenStudy (anonymous):

i didnt even notice you got it wrong

OpenStudy (anonymous):

its 8,

OpenStudy (anonymous):

how

OpenStudy (anonymous):

a + 2b - c 6+ 2*3-4 =8

OpenStudy (anonymous):

because its 6+2(3)-4

OpenStudy (anonymous):

which simplified is 8 as above^

OpenStudy (anonymous):

you simplified 6+2*3-4=8*3=24-4=20=2